# Improved Contact Graph Routing in Delay Tolerant Networks with Capacity and Buffer Constraints

### v1

@@ Abstract
Satellite communications present challenging characteristics. Continuous end-to-end connectivity may not be available due to the large distances between satellites. Moreover, resources such as link capacity and buffer memory may be limited. Routing in satellite networks is therefore both complex and crucial to avoid packet losses and long delays. The Delay Tolerant Network (DTN) paradigm has emerged as an efficient solution for managing these challenging networks. Contact Graph Routing (CGR), a deterministic routing algorithm, is one of the most popular DTN algorithms. CGR is compatible with the “store, carry, and forward” principle, whereby a node receives a message and stores it in its buffer until a transmission opportunity becomes available. However, CGR relies on simplified models to incorporate potential constraints in the route search. For instance, the linear volume assumption is often used to consider capacity constraints. Moreover, capacity management and buffer management are mostly performed during the forwarding phase, once an issue has occurred. In this paper, we propose to take measures before or during the route search in order to find routes that respect both contact-capacity limits and node-buffer limits. We introduce the contact splitting and edge pruning operations to effectively account for the routing constraints. This ensures that CGR outputs the optimal solution among the subset of valid solutions. The proposed approach can also be used to book resources to be used in case of issues during the forwarding step.
@@ I Introduction
Satellite communications belong to the category of Intermittently Connected Networks (ICNs). They are characterized by intermittent connectivity between the nodes of the network, preventing continuous end-to-end connections and resulting in high delays. This intermittent nature induces multiple discontinuous episodes of communication opportunities between pairs of nodes (so called contacts, see the following sections). A transmission can thus be delayed if there is no immediate communication opportunity. Nodes must therefore operate on a “store, carry, and forward” principle.
@@
As a result, since traditional Internet protocols are not suited to those challenges, the need for new protocols arose. The Delay Tolerant Network (DTN) [1] architecture with a new protocol layer has been proposed by the Internet Engineering Task Force (IETF) to handle these issues. The Bundle Protocol (BP) [2], considered by the Consultative Committee for Space Data Systems (CCSDS), is an instance of such protocols.
As mentioned above, routing is challenging with such discontinuous and time-dependent connections. The main routing issue in ICNs is the inability to establish a complete route due to lack of connectivity. Moreover, in addition to the delivery-time objective, optimizing delivery rate and resource consumption should also be considered. For instance, the use of transmission resources must be carefully optimized because of their limited availability. Additionally, as data units are often held in the node buffers for long periods, optimizing storage resources to avoid congestion and loss is crucial. Efficient routing algorithms are therefore needed in addition to suitable network capabilities via the above-mentioned protocols.
@@
Routing algorithms in DTNs can be classified into three categories depending on the availability of network information. They can be opportunistic [3], probabilistic [4] or deterministic, depending on whether the contacts are unpredictable, determined with a certain probability, or completely predetermined, respectively. In this paper, we focus on space communications, where natural and artificial satellites have deterministic motions. Thus, contacts can be predetermined, making deterministic routing algorithms suitable. The widely used deterministic routing algorithm in DTNs is the Contact Graph Routing (CGR) [5].
Since its invention, CGR has gained attention in the research community and several improvements have been proposed. For instance, using Dijkstra’s algorithm for the route search and choosing the Best-case Delivery Time (BDT) as objective were presented in [6]. Then, the Earliest Transmission Opportunity (ETO) was introduced to compute the time at which a bundle is transmitted through the contact [7]. The ETO takes into account the time a bundle has to wait before being transmitted because of other scheduled bundle transmissions on the same contact. Later, a complementary overbooking management was added to the ETO to improve routing decisions [8]. These enhancements among others result in the currently used version of CGR [5]. Nevertheless, this version provides only local and reactive contact-capacity and buffer-occupancy managements, which leads to sub-optimal solutions. Therefore, it is important to explore alternative approaches.
@@
One modern application, where capacity and buffer managements may be critical, is cislunar communications. With the advent of the Artemis Program [9], many projects in the lunar environment arose, triggering a high need of communication resources [10, 11]. Another application is low-earth orbit satellite constellations: These constellations are expected to be utilized extensively in the future. However, not all future constellations will be large enough to maintain end-to-end connectivity at all times. Even large constellations may not deploy all their satellites at the initial stage, requiring the establishment of a network with a limited number of satellites during the initial phase. As a result, many constellations are expected to temporarily buffer data.
@@
Main contributions. In this paper, we provide a routing solution based on CGR with a global and proactive congestion control. We propose to perform contact-capacity management and buffer-limit management before or during the route search such that the obtained route is optimal and respects these two constraints. The objective is achieved as follows.
@@
  • We propose to track node-buffer occupancy as function of time using forecast buffer tables. The buffer occupancy is then managed by:
@@
    • Temporarily forbidding some edges between contacts during the route search for a specific bundle to avoid buffer overflows.
The modified network information enables to remove the routes that do not respect the constraints (and only these routes). Subsequently, the route-search algorithm, such as CGR, can find the optimal solution among the subset of valid solutions.
The proposed approach can also be considered to book resources to be utilized in case of failure events during the forwarding phase. These resources, referred to as safety margin in the paper, are to be used by intermediate nodes having limited information on the network state.
Structure of the paper. The paper is structured as follows. In Section II, we first introduce key terms used in the paper and we provide a review of CGR. In Section III, we focus on existing capacity an buffer management challenges in DTNs to introduce our motivations. The proposed solution is introduced in Section IV. The benchmark source-routing CGR, considered as benchmark, is detailed in Section V. Simulation assumptions and results comparing our method with the benchmark are given in Sections VI and VII, respectively.
@@ II-A Key notions
  • Contact: A contact $C_{j}$ is defined by the communication opportunity between two nodes. It is thus characterized by sender and receiver nodes, a start time $t_{1,j}$, an end time $t_{2,j}$, and a data rate/capacity $R_{C_{j}}$ in bits per second (bps). The volume of a contact is its rate multiplied by its duration.
  • Contact Plan (CP): A CP is the list of contacts for a time period. It reflects the evolution of the network topology over time. An example is provided in Figure 2.
  • Edge: An edge exists between two contacts when the receiver of one contact matches the sender of a second contact, and the end time of the latter is later than the start time of the former.
  • Bundle: A bundle is the data unit encoded by the BP. A bundle $i$ is characterized by its size $S_{i}$ in bits. A bundle is generated at a source node and is provided with a destination node.
  • Route: A route or a path is a series of contacts that are used to deliver the bundle from the source node to the destination node.
Figure 1: Snapshot of a satellite communication scenario at time $t$ where $t_{1,1}\leq t\leq t_{2,1}$ and $t_{1,4}\leq t\leq t_{2,4}$ (see Figure 2).
II-B Contact graph routing
@@ II-B2 Route search
For a given destination, a route between the source node and the destination node is computed using Dijkstra’s algorithm [13]. However, an adaptation of this algorithm is used to consider the discontinuities of ICNs. The main idea is to use a contact-graph representation of the network rather than a node-graph representation. Dijkstra’s algorithm is thus implemented on the contact-graph representation of the network. The output of a Dijkstra search is the route from the source to the destination having the BDT.
@@
Yen’s algorithm [14] enables to determine a list of $k$ shortest routes by iteratively exploring the routes between a source and a destination, using Dijkstra with the contact-graph representation while avoiding loops and previously discovered paths.
@@
In the standard CGR implementation, each node computes a list of $k$ shortest routes to each node in the network using Yen’s algorithm, where $k$ is an input parameter. Then, once a bundle reaches a given intermediate node, the node determines a feasible precomputed route from its route list to determine the next hop.
II-B3 Validation and forwarding
@@
During the forwarding phase, the local queuing state of the buffer and the bundle properties are available. The remaining contact volume and the buffer queuing time for bundles scheduled on the same contact can then be taken into account. These checks, among others, help verify the validity of previously computed routes in the route list and prevent local congestion.
A list of candidate valid routes can thus be determined from the $k$ shortest routes. A valid route should guarantee contact volume availability given the size of the bundle. Moreover, one should also check that, taking into account transmission and queuing time, the bundle is effectively transmitted before the end of the corresponding contact, see ETO [7].
The best route among valid routes is selected to determine the bundle’s next hop. Once a route has been selected, the bundle is queued for transmission, and the available volume of the contact used is decreased by the size of the bundle.
II-B4 Additional comments on CGR

### v2

@@ Abstract
Satellite communications present challenging characteristics. Continuous end-to-end connectivity may not be available due to the large distances between satellites. Moreover, resources such as link capacity and buffer memory may be limited. Routing in satellite networks is therefore both complex and crucial to avoid packet losses and long delays. The Delay Tolerant Network (DTN) paradigm has emerged as an efficient solution for managing these challenging networks. Contact Graph Routing (CGR), a deterministic routing algorithm, is one of the most popular DTN algorithms. CGR is compatible with the “store, carry, and forward” principle, whereby a node receives a message and stores it in its buffer until a transmission opportunity becomes available. However, CGR relies on simplified models to incorporate potential constraints in the route search. For instance, the linear volume assumption is often used to consider capacity constraints. Moreover, capacity management and buffer management are mostly performed during the forwarding phase, once an issue has occurred. In this paper, we propose taking measures before or during the route search in order to find routes that respect both contact-capacity and node-buffer limits. We introduce the contact splitting and edge pruning operations to effectively account for the routing constraints. This ensures that CGR outputs the optimal solution among the subset of valid solutions. The proposed approach can also be used to book resources to be used in case of issues during the forwarding phase.
@@ I Introduction
Satellite networks are becoming increasingly important for many applications such as global coverage [1] and scientific missions in space [2]. These networks often face challenges such as long delays, limited connectivity, dynamic topology, and limited resources. Routing algorithms play an important role in this category of networks to ensure efficient data transmission.
As a first networking solution for satellite systems, traditional Internet Protocol (IP)-based routing algorithms have been adapted for satellite scenarios where the network topology can be assumed stable and where the connectivity can be considered continuous. These IP-based algorithms require explicit topology-update notifications [3][4]. In case of frequent network-topology changes, this approach therefore results in significant overhead which can drain the limited storage and processing resources of satellites. Moreover, these IP-based algorithms are designed for scenarios where end-to-end connectivity is available. However, satellite networks are often characterized by intermittent connectivity. This intermittent nature induces multiple discontinuous episodes of communication opportunities between pairs of nodes (so-called contacts, see the following sections). Limited effectiveness is therefore observed in non-dense satellite scenarios, such as the initial-stage deployment of a constellation.
@@
Since traditional Internet protocols are not suited to the mentioned challenges, the need for new protocols arose. The Delay Tolerant Network (DTN) [5] architecture with a new protocol layer has been proposed by the Internet Engineering Task Force (IETF) to handle these issues. The Bundle Protocol (BP) [6], considered by the Consultative Committee for Space Data Systems (CCSDS), is an instance of such protocols. One of the main advantages of a DTN architecture is the ability to “store, carry, and forward” data when immediate transmission is not possible. DTNs have become essential in modern space communications and are being used in recent projects [7, 8].
@@
In addition to suitable network capabilities via the above-mentioned protocols, efficient routing algorithms are also needed. Routing algorithms in DTNs can be classified into three categories depending on the availability of network information. They can be opportunistic [9, 10, 11], probabilistic [12], or deterministic [13], depending on whether the contacts are unpredictable, determined with a certain probability, or completely predetermined, respectively.
Opportunistic algorithms rely on message replication, aiming to ensure that at least one copy reaches the destination. They are often categorized as epidemic algorithms [10]. These algorithms are highly efficient when energy, buffer space, and link capacity are high or unlimited, as they achieve high delivery ratio and low latency. However, such ideal conditions are unrealistic, especially in satellite communication systems where capacity and buffer space are limited and must be used efficiently. To address this challenge, algorithms that limit the number of copies have been proposed [9, 14, 11]. In the latter reference, the number of copies generated by each node is adjusted based on its resource availability. Nevertheless, even with limited copies, resource wastage persists, as the forwarding process remains opportunistic. In satellite networks, this wastage can be significantly minimized by leveraging the predictable nature of satellite motion.
Probabilistic algorithms use statistical models and historical data to predict future network information (i.e., node encounters or contacts) [15]. However, such algorithms are not useful for most satellite networks where node connectivity patterns can be pre-determined.
Deterministic algorithms use comprehensive and pre-computed network information, such as satellite connectivity windows (i.e., contacts) for efficient routing. Unlike opportunistic methods, deterministic algorithms avoid unnecessary transmissions by calculating optimal routes in advance, ensuring that packets (also called bundles, the data units of the BP) are forwarded only when a path to the destination is available. This approach aligns well with the predictable topology of satellite networks, where orbital mechanics allow precise predictions of link availability.
In this paper, we focus on space communications, where natural and artificial satellites have deterministic motions. Thus, contacts can be predetermined, making deterministic routing algorithms suitable. The widely used deterministic routing algorithm in DTNs is Contact Graph Routing (CGR) [13]. CGR is well suited to satellite systems with limited resources. By prioritizing routing based on link availability and predetermined contact duration, the algorithm minimize unnecessary transmissions and optimize delivery time. The reader is invited to consult Section II-B for a review of recent improvements of CGR.
Moreover, in addition to the delivery-time objective, optimizing the delivery rate and the resource consumption should also be considered. For instance, the use of transmission resources must be carefully optimized because of their limited availability. Additionally, as data units are often held in the node buffers for long periods, optimizing storage resources to avoid congestion and loss is crucial. Nevertheless, the current version of CGR and the related latest research do not fully consider or fail to optimize contact capacity and buffer occupancy. Instead, they only provide local and reactive management during the forwarding phase. This leads to potential resource waste, sub-optimal solutions, and delays, especially in a resource-constrained environment. It is therefore important to explore alternative approaches.
@@
One modern application where capacity and buffer management may be critical is cislunar communications. With the advent of the Artemis Program [2], many projects in the lunar environment have arisen, triggering a high need for communication resources [16, 17]. Another application is low-earth orbit satellite constellations, which are expected to be utilized extensively in the future.
In this paper, we highlight the need to enhance the DTN routing algorithms, mainly CGR, to consider capacity and buffer management before or during the route search, bridging a critical gap in the current literature.
@@
Main contributions. We provide a routing solution based on CGR with a global and proactive congestion control. We propose to ensure contact-capacity management and buffer-limit management before or during the route search such that the obtained route is optimal and respects these two constraints. The objective is achieved as follows.
@@
  • We propose to track node buffer occupancy as function of time using forecast buffer tables. The buffer occupancy is then managed by:
@@
    • Temporarily forbidding some edges between contacts (i.e., a given contact succession) during the route search for a specific bundle to avoid buffer overflows.
The modified network information enables the removal of routes that do not respect the constraints (and only these routes). Subsequently, the route-search algorithm, such as CGR, can find the optimal solution among the subset of valid solutions.
The proposed approach can also be used to reserve resources to be utilized in case of failure events during the forwarding phase. These resources, referred to as safety margin in the paper, are to be used by intermediate nodes having limited information on the network state.
Structure of the paper. The paper is structured as follows. In Section II, we first introduce key terms used in the paper and we provide a review of CGR. In Section III, we focus on existing capacity and buffer management challenges in DTNs to introduce our motivations. The proposed solution is described in Section IV. The source-routing CGR benchmark is detailed in Section V. Simulation assumptions and results comparing our method with the benchmark are given in Sections VI and VII, respectively.
@@ II-A Key notions
  • Contact: A contact $\text{C}_{j}$ is defined by a communication opportunity between two nodes. It is characterized by sender and receiver nodes, a start time $t_{1,j}$, an end time $t_{2,j}$, and a data rate/capacity $R_{\text{C}_{j}}$ in bits per second (bps). The volume of a contact is its rate multiplied by its duration.
  • Contact Plan (CP): A CP is the list of contacts for a time period. It reflects the evolution of the network topology over time. An example is provided in Table I.
  • Edge: An edge exists between two contacts when the receiver of the first contact matches the sender of the second contact, and the end time of the first contact is earlier than the start time of the second.
  • Bundle: A bundle is the data unit encoded by the BP. A bundle $i$ is characterized by its size $S_{i}$ in bits. A bundle is generated at a source node and is assigned a destination node.
  • Route: A route or a path is a series of contacts that are used to deliver a bundle from its source node to its destination node.
Figure 1: Snapshot of a satellite communication scenario at time $t$ (e.g., $t_{1,1}\leq t\leq t_{2,1}$ and $t_{1,4}\leq t\leq t_{2,4}$, see Table I).
II-B Contact graph routing (history)
Since its invention, CGR has gained attention in the research community and several improvements have been proposed. Using Dijkstra’s algorithm for the route search and choosing the Best-case Delivery Time (BDT) as objective were presented in [18]. Then, the Earliest Transmission Opportunity (ETO) was introduced to compute the time at which a bundle is transmitted through the contact [19]. The ETO takes into account the time a bundle has to wait before being transmitted because of other scheduled bundle transmissions on the same contact. Later, a complementary overbooking management was added to the ETO to improve routing decisions [20]. Nevertheless, these managements are performed during the forwarding phase. Furthermore, a route list-management technique using Yen’s algorithm was proposed to further enhance the standard CGR algorithm [21]. Note however that this route management technique increases processing complexity during scheduling [22]. These enhancements among others result in the currently used version of CGR [13].
Research continues to improve the latest version of CGR, focusing on the optimization of limited resources including energy [23], computation [24, 25], buffer memory [26], and link capacity [27].
II-C Contact graph routing (implementation)
@@ II-C2 Route search
Route search is performed using an adapted version of Dijkstra’s algorithm [29]. This adaptation considers the discontinuities of Intermittently Connected Networks (ICNs) by using a contact-graph representation of the network rather than a node-graph representation. The output of a Dijkstra search is the route, consisting of a succession of contacts from the CP, connecting the source to the destination with the lowest BDT.
The detailed implementation can be found in [13]. Here, we explain the aspects of the route search that are essential for understanding our contribution. The idea is to explore the available contacts in the CP to determine the shortest route. The exploration of the CP is performed by analysing successively selected contacts denoted by $\text{C}_{s}$. For the first iteration, the selected contact $\text{C}_{s}$ is a root contact, which is virtual and whose sender and receiver are the bundle’s source node. At each iteration, a new contact $\text{C}_{s}$ is selected and used in the next iteration. Two main procedures are executed at each iteration. The first procedure is the Contact Review Procedure (CRP), used to explore the CP. The second one is the Contact Selection Procedure (CSP), used to determine the following $\text{C}_{s}$.
Algorithm 1 presents both the original CRP and our contributions. Our modifications are marked with asterisks (**, lines $2$ and $8-12$) in the algorithm. In this section, we explain only the original CRP. The modifications will be explained in Section IV-B2.
In the CRP, the CP is explored starting from a selected contact $\text{C}_{s}$. During this procedure, all contacts are considered for exploration. In the algorithm, the following characteristics of a contact are updated and used:
  • A successor of a contact. I.e., the conctact that comes after in a route.
  • A predecessor of a contact. I.e., the contact that comes before in a route.
  • An arrival time $t_{\text{arr},i}$ that indicates the time needed to reach the receiver node of the contact.
  • Visited nodes, which are the nodes that exist in the route or the succession of contacts that led to this contact.
  • A visited flag, which indicates if this contact has been visited, meaning that it has been previously selected and used to start exploring contacts in the CP.
  • A suppressed indicator, which indicates if the use of the contact is forbidden (for security reasons for example).
A contact $\text{C}_{i}\in$ CP must meet the following criteria to be explored (otherwise it is ignored) (lines $5-7$). If these criteria are met $C_{i}$ will be considered as a potential successor of $C_{s}$.
a) $\text{C}_{i}$ must have the selected contact’s receiver as its sender.
b) Let $t_{\text{arr},i}$ be the arrival time at $\text{C}_{i}$, i.e., the time at which the receiver node of contact $\text{C}_{i}$ is reached through the succession of contacts that led to $\text{C}_{i}$. The end time of $\text{C}_{i}$ must be greater than the arrival time at the receiver of the selected contact, $t_{\text{arr},s}$. This means that the successor $\text{C}_{i}$ must not end before the bundle arrives at its sender node (the receiver node of $\text{C}_{s}$).
c) $\text{C}_{i}$ should not have been visited, meaning that it should not have been selected as $\text{C}_{s}$ in a previous CRP iteration.
d) The receiver node of $\text{C}_{i}$ must not be visited i.e., be a node of the contact sequence that led to $\text{C}_{i}$, to avoid loops.
e) $\text{C}_{i}$ must not be suppressed (e.g., removed or forbidden).
f) The remaining volume of $\text{C}_{i}$ must be greater than zero. This means that after allocating $\text{C}_{i}$ to bundles and deducting bundle size from the volume of $\text{C}_{i}$, some volume must remain available in $\text{C}_{i}$ to allow more data to pass through.
Once a contact has passed all checks, it is considered as an explored contact. The arrival time of an explored contact is computed according to the time taken to reach the receiver of the explored contact following the route that led to it (line $13$). If the computed arrival time is less than its last calculated value $t_{\text{arr},i}$ (in the case where $\text{C}_{i}$ has already been explored), it is updated (lines $14$ and $15$). In addition, the selected contact $\text{C}_{s}$ becomes the predecessor of $\text{C}_{i}$ and the visited nodes of $\text{C}_{i}$ are updated by adding the visited nodes of the nodes within the succession of contacts that led to $\text{C}_{i}$ (lines $16$ and $17$) (to avoid getting back to them, which would induce loops).
Each time the destination node is reached (i.e., when the receiver of an explored $C_{i}$ is the bundle’s destination), the arrival time is saved as BDT if it is lower than the previously calculated BDT. In such cases, $\text{C}_{i}$ is considered as a final contact $\text{C}_{\text{end}}$ with its corresponding BDT (lines $18-21$). After finishing all contacts in the CP, the currently selected contact $\text{C}_{s}$ is marked as visited to avoid repeating the same work later (line $24$).
Algorithm 1 Updated Contact Review Procedure (CRP)
1:Input: CP, $\text{C}_{s}$, $\text{C}_{\text{end}}$, BDT
2:** Additional input: Bundle b, $T_{1,s}=\{t_{1,o_{l}}\}\mid o_{l}\in\mathcal{O}_{Z_{1}}$, $T_{2,s}=\{t_{2,o_{l}}\}\mid o_{l}\in\mathcal{O}_{Z_{1}}$, where ${Z_{1}}$ is the receiver of $\text{C}_{s}$*
3:Output: updated CP, $\text{C}_{\text{end}}$, BDT
4:for $\text{C}_{i}\in$ CP do
5:     if sender of $\text{C}_{i}\neq$ receiver of $\text{C}_{s}$ or $t_{i,2}\leq t_{\text{arr},s}$ or $\text{C}_{i}$ is already visited or receiver of $\text{C}_{i}$ is a visited node by $\text{C}_{s}$ or $\text{C}_{i}$ is suppressed or remaining volume of $\text{C}_{i}==0$ then
6:         ignore $\text{C}_{i}$
7:     end if
8:** Additional condition
9:     $T_{1,s}=[T_{1,s},\infty]$
10:     $T_{2,s}=[0,T_{2,s}]$
11:     if any ($t_{\text{arr},s}\geq T_{2,s}$ and $t_{i,1}\geq T_{1,s}$then
12:         ignore $\text{C}_{i}$
13:     end if
14:**
15:     compute $t_{\text{arr}}$ to the receiver of $\text{C}_{i}$, which follows $\text{C}_{s}$
16:     if $t_{\text{arr}}<t_{\text{arr},i}$ then
17:         $t_{\text{arr},i}=t_{\text{arr}}$
18:         predecessor of $\text{C}_{i}=\text{C}_{s}$
19:         $\text{C}_{i}$’s visited nodes $=\text{C}_{s}$’s visited nodes and $\text{C}_{i}$’s     receiver
20:         if $\text{C}_{i}$’s receiver is bundle’s destination and     $t_{\text{arr},i}<$ BDT then
21:              BDT $=t_{\text{arr},i}$
22:              $\text{C}_{\text{end}}=\text{C}_{i}$
23:         end if
24:     end if
25:end for
26:Consider $\text{C}_{s}$ as visited
In the CSP, the best contact among the explored contacts from CRP is selected based on the shortest arrival time. This contact becomes the new selected contact for the next CRP iteration. The iterations continue until no additional contacts can be selected for further exploration. The final contact of the route is then the one with the lowest BDT.
Once these procedures are complete, the route is constructed in reverse order. Starting from the final contact (which has the lowest BDT), the algorithm traces back through the predecessors until reaching the root contact. For each Dijkstra call, the CP is reset to its initial state (i.e., as if no contacts had been explored).
@@
Yen’s algorithm [30] enables the determination of the $k$ shortest routes by iteratively exploring paths between a source and a destination. It uses Dijkstra’s algorithm with the contact-graph representation while avoiding loops and previously discovered paths.
@@
In the standard CGR implementation, each node computes a list of the $k$ shortest routes to each node in the network using Yen’s algorithm, where $k$ is an input parameter. This is referred to as per-hop routing. When a bundle reaches an intermediate node, this node selects a feasible pre-computed route from its route list to determine the next hop.
II-C3 Validation and forwarding
@@
In the forwarding phase, the system examines both the remaining contact volume and the buffer queuing time for bundles scheduled on the same contact. These validation checks verify the feasibility of previously computed routes and prevent local congestion.
The system determines valid route candidates from the $k$ shortest routes. A valid route must guarantee sufficient contact volume for the bundle’s size. Additionally, the system must verify that the bundle can complete transmission, including queuing time, before the contact ends (see ETO [19]).
The system selects the optimal route from valid candidates to determine the bundle’s next hop. After the route selection, the system queues the bundle for transmission and decreases the contact’s available volume by the bundle’s size.
II-C4 Additional comments on CGR